please help me with this problem Thanks
It said solve the linear equation using the substitution method


5x-2y=6
y=2x-3

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

5x-2y=6
y=2x-3
--------------------
Substitute 2nd into the 1st to get:
5x -2(2x-3) = 6
5x -4x + 6 = 6
x = 0
----------
Substitute that into y = 2x-3 to solve for "y":
y = 2*0 - 3
y = -3
==============
Cheers,
Stan H.

Answer by Alan3354(69443)

5x-2y=6
y=2x-3
Since y is already expressed as a function of x, sub for y in the 1st eqn.
---------
5x -2(2x-3) = 6
5x - 4x + 6 = 6
5x = 4x
There is no solution. This is similar to stating x+1 = 3, then x+1 = 4. No solution is possible.